# cammarl: Conformal Action Modeling in Multi Agent Reinforcement Learning

### v1

@@ 5 Experiments and Results
Advancing from the inputs available in NOAM, we implement TAAM by allowing $\mathcal{N}_{self}$ to additionally utilize $\mathcal{N}_{other}$’s true actions to train. This baseline helps us evaluate cammarl against works that estimate other agents’ actions in the environment and use those predictions to enhance the decision-making of their controlled autonomous agents. By giving the true actions as inputs, this baseline can act as an upper bound to such works (for example, [16, 12, 48, 26, 32, 1]). Figure 4 shows TAAM’s performance curve. Using additional information, it does reasonably better than NOAM in both tasks.
@@
As discussed in Section 2, learning world models often involves reconstructing observation as an additional task while learning task-related policies [18, 14, 15]. Inspired by this research, we implement the TOAM baseline where we allow access to $\mathcal{N}_{other}$’s true observations to $\mathcal{N}_{self}$ during training and execution. In other words, we augment $\mathcal{N}_{self}$’s partial observations with the other agent’s local observations too. This baseline can act as an upper bound to the performances of research works that learn to reconstruct states for agents[14, 15, 44, 36, 46]. As expected, $\mathcal{N}_{self}$ performs considerably better than NOAM in both environments (Figure 4). Here, the difference in returns in TOAM and TAAM can be attributed to the fact that the local observations of other agents include more information that can be useful to infer their behavior. For instance, in CN, knowing the relative positions of other agents with respect to the landmarks can be more useful to infer which landmark that agent might be approaching when compared to knowing its current (or history) actions.
@@
Building over the inputs of TOAM, we construct another but stronger baseline, EAP, in which $\mathcal{N}_{self}$ uses an additional neural network classifier to model a probability distribution over $\mathcal{N}_{other}$’s actions. In other words, instead of predicting conformal sets of actions (like in cammarl), in this baseline, $\mathcal{N}_{self}$ tries to model $\mathcal{N}_{other}$’s actions from the latter’s observations without accounting for any uncertainty quantification. This baseline is inspired by works that explicitly model the other agent’s actions in the environments and utilize them to inform their controlled agent’s decision-making (for instance, [16, 12, 48]). Hence, here, a cross-entropy loss is used to train the added sub-module that predicts the $\mathcal{N}_{other}$’s actions along with a PPO loss to train $\mathcal{N}_{self}$’s policy network. Figure 4 shows that cammarl agents are able to distinctly perform better than EAP in LBF, however, interestingly, the performance curve for this baseline nearly overlaps with cammarl in CN. Also, in LBF, the curves for TOAM and EAP seem to significantly overlap. We speculate that in a complicated task like LBF, estimating the exact action of $\mathcal{N}_{other}$ can be difficult, and with unaccounted uncertainty in the predictions, $\mathcal{N}_{self}$ suffers from a lower return. In CN, which is comparatively simpler, the closeness of returns in EAP and cammarl seem reasonable as even the conformal model predictions eventually start predicting the most probable action with higher probabilities and hence a set of size one (more on this in Section 6).
@@
Now, we implement cammarl, where the conformal action prediction model periodically trains on collected observations of $\mathcal{N}_{other}$ and predicts a corresponding conformal set of actions. $\mathcal{N}_{self}$ uses these estimates of $\mathcal{N}_{other}$’s actions along with its own observations and then decides upon its actions in the environment. Figure 4 shows that cammarl agents obtain returns that are much closer to the upper bound, GIAM, than the lower bound, NOAM. Furthermore, cammarl’s better performance compared to TOAM in both environments can be attributed to the fact that it can be difficult to predict the $\mathcal{N}_{other}$’s intentions by only using $o_{other}$ without any information pertaining to its actions in those situations. And, in TAAM, $\mathcal{N}_{self}$ is expected to implicitly encode information regarding $\mathcal{N}_{other}$’s observations from its own local observations or in the latent space and map it to $\mathcal{N}_{other}$’s true actions. We speculate that this could be a strong assumption and consequently very difficult, hence, cammarl agents outperform TAAM too. Note here that the sets output by the conformal action prediction model are of varying sizes in each iteration. Now, to be able to use these dynamically changing inputs for $\mathcal{N}_{self}$ in cammarl, we convert the output sets to a corresponding binary encoding (by firing up the bits in a zero vector at indices corresponding to the actions predicted by the model). We discuss some more ways to be able to use conformal prediction sets with dynamic sizes and compare cammarl’s performances in all variations later in Appendix A.
@@
In summary, through experiments in two complex cooperative tasks, we show that (1) cammarl indeed works, (2) it outperforms common settings like NOAM, TOAM, and TAAM which assume the availability of other agents’ true trajectories during training and execution (generally infeasible in real-world scenarios), (3) Its performance is closest to our upper bound of performance (GIAM), (4) cammarl agents learn their policies faster than the other baselines, and (5) cammarl can be preferred over strong benchmarks such as EAP owing to its higher interpretability due to the theoretical guarantees of conformal predictions in terms of coverage [4] (discussed more in Section 6).
@@ 6 Discussion
We collected the set sizes produced in cammarl throughout the training and report them in Figure 4(a) and  4(e). Smaller sets are preferred, as they carry specific information which can be more useful practically. The curves show a decreasing trend in the set sizes in cammarl in both CN and LBF respectively when tracked over the number of updates of the conformal prediction model during training. This is a good sign for cammarl, as it shows that the conformal predictions are becoming more precise with continued training over time.
@@
As also discussed earlier, it is desirable for the predicted sets to provide 1 - $\alpha$ coverage for a pre-defined user-specified $\alpha$ such as 10%. Formally, to map a feature vector, $o_{other}\in\mathcal{O}_{other}$, to a subset of discrete responses, $a^{\prime}_{other}\in\mathcal{A}_{other}$, it is useful to define an uncertainty set function, $\mathcal{C}$($o_{other}$), such that $P(a^{\prime}_{other}\in C(o_{other}))\geq 1-\alpha$. Figure 4(b) and  4(f) shows the increasing trend of confidence coverage in cammarl.
@@ 7 Conclusion
In this article, we propose a novel MARL algorithm, cammarl, which calls for confident reasoning about other artificial agents in the environment and benefiting from inferences about their behavior. Through experiments in two cooperative multi-agent tasks, CN and LBF, we showed that guiding an agent’s decision-making by inferring other agents’ actions in the form of conformal sets, indeed helps in achieving better performances of the learning agents. By using conformal prediction, we were also able to ensure the estimation of predictive sets that covered the real predictions of the intentions of other agents with a very high pre-specified probability of 95%. There could be several future directions from here. First, cammarl is certainly generalizable to bigger networks or more simple classifiers, and also analyzing its changing performance on varying buffer sizes can help in better comprehending its efficiency. Second, it would be interesting to investigate the cammarl’s scalability to a system of many agents (say 100 or 1000) or on more complicated multi-agent environments such as tasks requiring a higher need for coordination. Finally, in this work, we restricted the agents to infer the behavior of other agents only via conformal sets; it would be interesting to study the cases where more ways of sharing information or modeling agents’ behavior are additionally allowed.
Acknowledgements.
The authors would like to thank the Digital Research Alliance of Canada for compute resources and CIFAR for research funding. The authors are also grateful for the detailed comments from Somjit Nath and all reviewers.
@@ Acknowledgements.
Appendix A Variations of the cammarl algorithm
@@
Appendix B MARL agents: Implementation details
We use proximal policy optimization (PPO) [40] to update the decision-making policy for both the RL agents, however, any other RL algorithm could be used alternatively. For the individual actor and critic networks, we used 2 fully-connected multi-layer perceptron (MLP) layers. For the conformal prediction model $\mathcal{C}$, we use another fully-connected MLP with 2 layers each with 64 hidden nodes. All the agents and the model $\mathcal{C}$ collect their individual experiences and train their own policies independently.
Appendix C Partially observable MDP

### v2

@@ 4.2 Results
Advancing from the inputs available in NOAM, we implement TAAM by allowing $\mathcal{N}_{self}$ to additionally utilize $\mathcal{N}_{other}$’s true actions to train. This baseline helps us evaluate cammarl against works that estimate other agents’ actions in the environment and use those predictions to enhance the decision-making of their controlled autonomous agents. By giving the true actions as inputs, this baseline can act as an upper bound to such works(He et al., 2016; Grover et al., 2018; Zintgraf et al., 2021; Mealing & Shapiro, 2015; Panella & Gmytrasiewicz, 2017; Albrecht & Ramamoorthy, 2015). Figure 3 shows TAAM’s performance curve. Using additional information, it does reasonably better than NOAM in both tasks.
@@
As discussed in Section 2, learning world models often involves reconstructing observation as an additional task while learning task-related policies (Jain et al., 2022; Hafner et al., 2020, 2021). Inspired by this research, we implement the TOAM baseline where we allow access to $\mathcal{N}_{other}$’s true observations to $\mathcal{N}_{self}$ during training and execution. In other words, we augment $\mathcal{N}_{self}$’s partial observations with the other agent’s local observations too. This baseline can act as an upper bound to the performances of research works that learn to reconstruct states for agents (Hafner et al., 2020, 2021; Wang et al., 2022; Park et al., 2019b; Zhang et al., 2021). As expected, $\mathcal{N}_{self}$ performs considerably better than NOAM in both environments (Figure 3). Here, the difference in returns in TOAM and TAAM can be attributed to the fact that the local observations of other agents include more information that can be useful to infer their behavior. For instance, in CN, knowing the relative positions of other agents with respect to the landmarks can be more useful to infer which landmark that agent might be approaching when compared to knowing its current (or history) actions.
@@
Building over the inputs of TOAM, we construct a stronger baseline, EAP, in which $\mathcal{N}_{self}$ uses an additional neural network classifier to model a probability distribution over $\mathcal{N}_{other}$’s actions. In other words, instead of predicting conformal sets of actions (like in cammarl), in this baseline, $\mathcal{N}_{self}$ tries to model $\mathcal{N}_{other}$’s actions from the latter’s observations without accounting for any uncertainty quantification. This baseline is inspired by works that explicitly model the other agent’s actions in the environments and utilize them to inform their controlled agent’s decision-making (for instance, (He et al., 2016; Grover et al., 2018; Zintgraf et al., 2021)). Hence, here, a cross-entropy loss is used to train the added sub-module that predicts the $\mathcal{N}_{other}$’s actions along with a PPO loss to train $\mathcal{N}_{self}$’s policy network. Figure 3 shows that cammarl agents are able to distinctly perform better than EAP in LBF, however, interestingly, the performance curve for this baseline nearly overlaps with cammarl in CN. Also, in LBF, the curves for TOAM and EAP seem to significantly overlap. We speculate that in a complicated task like LBF, estimating the exact action of $\mathcal{N}_{other}$ can be difficult, and with unaccounted uncertainty in the predictions, $\mathcal{N}_{self}$ suffers from a lower return. In CN, which is comparatively simpler, the closeness of returns in EAP and cammarl seem reasonable as even the conformal model predictions eventually start predicting the most probable action with higher probabilities and hence a set of size one (more on this in Section 6).
@@
Now, we implement cammarl, where the conformal action prediction model periodically trains on collected observations of $\mathcal{N}_{other}$ and predicts a corresponding conformal set of actions. $\mathcal{N}_{self}$ uses these estimates of $\mathcal{N}_{other}$’s actions along with its own observations and then decides upon its actions in the environment. Figure 3 shows that cammarl agents obtain returns that are much closer to the upper bound, GIAM, than the lower bound, NOAM. Furthermore, cammarl’s better performance compared to TOAM in both environments can be attributed to the fact that it can be difficult to predict the $\mathcal{N}_{other}$’s intentions by only using $o_{other}$ without any information pertaining to its actions in those situations. And, in TAAM, $\mathcal{N}_{self}$ is expected to implicitly encode information regarding $\mathcal{N}_{other}$’s observations from its own local observations or in the latent space and map it to $\mathcal{N}_{other}$’s true actions. We speculate that this could be a strong assumption and consequently very difficult, hence, cammarl agents outperform TAAM too. Note here that the sets output by the conformal action prediction model are of varying sizes in each iteration. Now, to be able to use these dynamically changing inputs for $\mathcal{N}_{self}$ in cammarl, we convert the output sets to a corresponding binary encoding (by firing up the bits in a zero vector at indices corresponding to the actions predicted by the model). We discuss some more ways to be able to use conformal prediction sets with dynamic sizes and compare cammarl’s performances in all variations later in the supplementary.
@@
In summary, through experiments in two complex cooperative tasks, we show that (1) cammarl indeed works, (2) it outperforms common settings like NOAM, TOAM, and TAAM which assume the availability of other agents’ true trajectories during training and execution (generally infeasible in real-world scenarios), (3) Its performance is closest to our upper bound of performance (GIAM), (4) cammarl agents learn their policies faster than the other baselines, and (5) cammarl can be preferred over strong benchmarks such as EAP owing to its higher interpretability due to the theoretical guarantees of conformal predictions in terms of coverage (Angelopoulos et al., 2020) (discussed more in Section 6).
4.3 Results with more than 2 agents
Figure 4: Comparison of agent performances (in terms of reward accumulation) in environments with more than 2 agents: (a) Pressure Plate and (b) Google Football. Interestingly, Pressure Plate cammarl can be seen to learn arguably faster, but all methods converge to similar results, whereas in Google Football, cammarl reaches a higher reward than the baselines. In Google Football, the observations are global, so we did not include TOAM and GIAM. The curves are averaged over five independent runs.
cammarl can be extended to problems with more than 2 agents, where the main agent uses a conformal model for each of the other agents. From Figure 4 (a) we see that although all the algorithms ultimately converge to the optimal reward, cammarl is more sample efficient and converges in roughly 10% fewer episodes compared to GIAM and TOAM. In Google Football (Figure 4 (b)) we notice a considerable improvement as well, with conformal predictions helping in much better reward accumulation. These experiments also highlight the benefit of using conformal predictions, as TAAM with accurate actions do not converge as fast as cammarl.
Figure 5: Comparison of agent performances (in terms of reward accumulation) in LBF with 2 agents (2p) and 6 foods (6f) respectively with cooperation turned off. cammarl’s performance is very close to the upper bound, GIAM, and is considerably better than the other extreme, NOAM. Interestingly, cammarl also seems to converge faster than the other baselines. The curves are averaged over five seeds.
Figure 6: Comparison of agent performances in CN with uncertainty information available to $\mathcal{N}_{self}$ during training. This graph highlights the merit of using Conformal Action Predictions over simple uncertainty estimation.
4.4 Results in Mixed Settings
For these experiments, we use the original LBF environment, but we turn off the cooperation flag. This is a slightly easier task, however, each agent here would fight for different food positions, so it is a mixed setting between cooperation and opposition. Modeling actions of other agents can also help to improve performance in this setting, although it is not as critical as in the cooperative setting.
Figure 5 highlights the training performance on LBF with 2 agents respectively. Although this is a simpler task, in which all algorithms roughly converge to the same reward, cammarl seems to be on par with our upper bound GIAM, especially in terms of faster convergence.
5 Motivating Conformal Predictions
In Figure 3 (a) and (b), we observe the improved performance of cammarl by predicting conformal sets. One key ingredient to cammarl was the addition of uncertainty predictions to the actions of the other agents. Thus, we can attribute the increased performance of cammarl to this as well. To test this theory, we added one more baseline where similar to the action prediction baseline (EAP), we add the action prediction probabilities directly into the state space. We call this Action Prediction with Uncertainty (APU), and both APU and cammarl operate on the same information. From Figure 6, we can conclude that just adding uncertain predictions is not enough to achieve an uplift in performance that we see for cammarl and there is definite merit to using conformal predictions. Another reason for the poor performance of APU would be that the agent has to parse through the relevant information and learn from it, whereas they are readily provided in a concise manner for cammarl.
@@ 6 Discussion
We collected the set sizes produced in cammarl throughout the training and report them in Figure 7 (a) and  7 (f). Smaller sets are preferred, as they carry specific information which can be more useful practically. The curves show a decreasing trend in the set sizes in cammarl in both CN and LBF respectively when tracked over the number of updates of the conformal prediction model during training. This is a good sign for cammarl, as it shows that the conformal predictions are becoming more precise with continued training over time.
@@
As also discussed earlier, it is desirable for the predicted sets to provide 1 - $\alpha$ coverage for a pre-defined user-specified $\alpha$ such as 10%. Formally, to map a feature vector, $o_{other}\in\mathcal{O}_{other}$, to a subset of discrete responses, $a^{\prime}_{other}\in\mathcal{A}_{other}$, it is useful to define an uncertainty set function, $\mathcal{C}$($o_{other}$), such that $P(a^{\prime}_{other}\in C(o_{other}))\geq 1-\alpha$. Figure 7 (b) and  7 (f) shows the increasing trend of confidence coverage in cammarl.
@@ 7 Conclusion
In this article, we propose a novel MARL algorithm, cammarl, which calls for confident reasoning about other artificial agents in the environment and benefiting from inferences about their behavior. Through experiments in two cooperative multi-agent tasks, CN and LBF, we showed that guiding an agent’s decision-making by inferring other agents’ actions in the form of conformal sets, indeed helps in achieving better performances of the learning agents. By using conformal prediction, we were also able to ensure the estimation of predictive sets that covered the real predictions of the intentions of other agents with a very high pre-specified probability of 95%.
Limitations and Future Works: In our paper, we analyzed cammarl with two agents, however, cammarl is certainly generalizable to bigger networks or more simple classifiers, and analyzing its changing performance on varying buffer sizes can help in better comprehending its efficiency. Second, it would be interesting to investigate the cammarl’s scalability to a system of many agents (say 100 or 1000) or on more complicated multi-agent environments such as tasks requiring a higher need for coordination. Thirdly, our mathematical model in Section 3.1 makes an assumption that the state space is accessible globally which may not be the case in some problems. Finally, in this work, we restricted the agents to infer the behavior of other agents only via conformal sets; it would be interesting to study the cases where more ways of sharing information or modeling agents’ behavior are additionally allowed.
@@
Appendix A Environments
Figure 8: Multi-agent cooperative environments used in this study: (a) OpenAI MPE’s Cooperative Navigation: Agents (blue) learn to cover the two landmarks (black) avoiding collisions. The figure shows cooperative navigation with 2 agents (N=2) and 2 landmarks (L=2). and (b) Level-based foraging: Agents must collect food and learn to cooperate using sparse rewards. This is a 12 $\times$ 12 level-based foraging grid-world with 2 cooperative players and 4 food locations. (c) Pressure Plate: Agents must stand on the pressure plate to keep the gates open for one of the agents to reach the goal. This is a pressure plate environment with 4 agents.
Cooperative Navigation (CN(Mordatch & Abbeel, 2017; Lowe et al., 2017).
where $|\mathcal{N}|$ is the number of agents and L are the landmarks in the environment. Here, c is the number of collisions in an episode and the agents are penalized with -1 for each time two agents collide.
Level-based foraging (LBF(Albrecht & Ramamoorthy, 2015).
In this environment, $\mathcal{N}_{self}$ and $\mathcal{N}_{other}$ are part of a 12$\times$12 grid world which contains four randomly scattered food locations, each assigned a level. The agents also have a level of their own. They attempt to collect food which is successful only if the sum of the levels of the agents involved in loading is greater than or equal to the level of the food. This is a challenging environment, requiring agents to learn to trade off between collecting food for their own and cooperating with the other agent to acquire higher team rewards. Moreover, this environment has sparse rewards making it difficult to learn and operate independently in the environment. In particular, each agent is rewarded equal to the level of the food it managed to collect, divided by its level (its contribution).
Pressure Plate [^3] https://github.com/uoe-agents/pressureplate
In this domain, the agents are expected to learn to co-operate in traversing the gridworld to collect the yellow treasure chest. Each agent is assigned a pressure plate that only they can activate. That agents must stand on the pressure plate to keep the corresponding doorway to open while the other agents can go through it leaving the agent behind. Each agent can only view a limited portion of the environment. If an agent is in the room with their assigned plate, their reward is the negative normalized Manhattan distance between the agent and plate’s position, else the reward is the difference between the current and desired room number.
Google Football (Kurach et al., 2020)
For our task, we use the 3 vs 1 with keeper game which is one of the 5 domains in the Football Academy of Google Research Football domain. The main agent controls 3 agents whose objective is to score a goal. The opponent, consisting of 1 defender and a keeper will try to prevent scoring the goal. This environment requires cooperation among the 3 players to score, including skills like passing the ball. The agents receive a reward when it successfully scores a goal past the opponent.
Appendix B Variations of the cammarl algorithm
@@
Appendix C MARL agents: Implementation details
C.1 Partially observable MDP
@@ C Conformal Prediction model
We use proximal policy optimization (PPO) (Schulman et al., 2017) to update the decision-making policy for both the RL agents, however, any other RL algorithm could be used alternatively. For the individual actor and critic networks, we used 2 fully-connected multi-layer perceptron (MLP) layers. For the conformal prediction model $\mathcal{C}$, we use another fully-connected MLP with 2 layers each with 64 hidden nodes. All the agents and the model $\mathcal{C}$ collect their individual experiences and train their own policies independently.
We add regularization to the conformal model to encourage small set sizes. This regularization is controlled by two hyper-parameters $\lambda$ and $k_{reg}$. Each time the conformal model is trained, we choose the $\lambda$ and $k_{reg}$ parameters from a set that optimizes for small sizes of the action set. The $k_{reg}$ values are dependent on the logits, however, the $\lambda$ values are in a set $[0.001,0.01,0.1,0.2,0.5]$.
Appendix D Visualization of Learned Policy
In this section, we visualize the learned policy in Pressure Plate with cammarl demonstrating cooperation between the 4 agents.
Figure 10: Visualization of the policy learned by cammarl in Pressure Plate with 4 agents